### v1

@@ 6 Results
In this section we present the empirical estimates of the impact of Cohesion Fund payments on regional economic performance. [^10] The estimation is conducted by using the gsynth library (Xu and Liu, 2021) in the statistical software R (R Core Team, 2021) The section is organized as follows: firstly, in section 6.1 we present the average treatment effect on the treated. Next, in section 6.2, we delve into the heterogeneous nature of the effects on the level of regional GVA per capita, focusing both on the sub-national (regional) and the national level. Then, in section 6.2.3 we turn our attention on the growth effects of the Cohesion Fund expenditures. Lastly, section 6.3 provides some graphical evidence on the nonlinear nature of the relationship between funding intensity and its GVA effects.
@@ 6.1 Average Treatment Effect
We start by presenting the average treatment effect across all regions as estimated in section 5.2.1; specifically, see equation (4). Figures 3 and  3 present the average treatment effect on the y-axis against years relative to treatment for GVA per capita and Gross Fixed Capital Formation (GFCF), respectively. The vertical red line that crosses the x-axis in year 0 represents the initial year in which each region received CF payments. Note that this year might differ for each region as EU countries became eligible for the CF payments at different years and, as such, the regions in our sample received payments at different points in time. Thus, the treatment effect is shown by the black line for each year that follows the zero year. The grey shaded area represents the upper and lower bound of the $95\%$ confidence interval calculated using a parametric bootstrap technique as in Xu (2017) (see Appendix A for more details).
@@
Figures 3 and 3 show that on average CF payments have a persistent positive effect on both the GVA and GFCF of a region that receives them (treated) compared to the counterfactual scenario where these funds are not provided by the EU. [^11] Figure C.1 in Appendix C plots both the observed and the respective estimated counterfactual average GVA in logs as well as in growth rates. Their difference is the average treatment effect presented in 3. Additionally, Figure C.2 depicts the same series for the growth rate of GVA. Our methodology allows us to provide a quantification of the average GVA effect in the regions which were eligible for Cohesion Fund payments. Specifically, on the seventh year under treatment the average treatment effect is equal to 0.086 (in logs), while for the entire period it averages to 0.05. Similarly, for the 14th year under treatment, the average treatment effect is equal to 0.091, while for the second 7-year period it is equal to 0.097 on average. This implies that, for the hypothetical EU27 average region, GVA per capita increases by €468 more than it would under the counterfactual scenario (i.e., without CF support). [^12] In this hypothetical EU27 average region, GVA per capita rises by €2,168 over the first seven years of treatment—from €9,064 in the first year to €11,232 in year seven. Under the counterfactual scenario, the corresponding increase is €1,700 —from €8,928 to €10,628 in year seven. After 14 years the average treatment effect is equal to 0.075 in logs. This implies that the hypothetical EU27 average region, GVA per capita increases by €330 more than it would under the counterfactual scenario. Following similar calculation, the hypothetical EU27 average region, GFCF per capita increases by €385 and €415 more than it would under the counterfactual scenario after 7 and 14 years under treatment, respectively. Lastly, as can be seen from the Figures, the majority of this rise occurs in the first seven years, while further gains beyond that period are relatively smaller. This is a first indication that the impact of the Cohesion Fund is, on average, frontloaded.
@@
Note: Average Treatment Effect on the Treated (ATT) over time (black solid line) and $95\%$ credible intervals (gray shaded area) for the log of GVA per capita (left panel) and the log of GFCF (right panel).
@@ 6.2.1 Quantitative analysis of the heterogeneous effects
The previous section focused on the average treatment effect across all regions that have received CF payments. However, a key advantage of our empirical strategy is that it allows us to explore the heterogeneous effects of CF across the different regions in our sample. Specifically, Figure 4 presents the entire distribution of the GVA effects across treated regions at 7 and 14 years after receiving CF for the first time along with the average effect across all years. Similarly, Figure 5 displays the analogous distributions for the GFCF effect.
A visual inspection of Figures 4 and  5 shows significant heterogeneity in the quantitative effect of CF in EU regions. This finding is novel, as most existing papers usually limit their analysis to the estimation of an average effect across regions, ignoring any potential heterogeneity. Both Figures suggest that the heterogeneous impacts should be closely accounted for and studied, as they can provide valuable information to EU policymakers regarding the optimal allocation of EU resources. Indicatively, the observed heterogeneity could be related to the intensity of the treatment, i.e. the size of the Cohesion Fund expenditures relative to GVA. If a maximum desired intensity is identified, this provides flexibility to re-allocate the resources and achieve better outcomes. We touch upon this issue in more detail in secton 6.3.
@@
Table 4: Summary statistics for the distribution of the regional effect on GVA per capita
7 years 14 years Average
[…]
Skewness 0.81 -0.06 0.15
Kurtosis 3.81 2.33 2.42
Table 5 presents summary statistics for the distribution of CF effects on gross fixed capital formation (GFCF). Although most of the patterns observed for GVA effects hold, some interesting differences emerge. Specifically, the impact of the CF on regional investment is significantly larger in magnitude, affecting a higher percentage of treated regions. Almost 90% of the treated regions experience a positive effect on their investment within the first 7 years compared to 64.5% for GVA. This proportion decreases over time, and 62. 5% of the treated regions still have a positive impact after 14 years. Both skewness and kurtosis are higher than the respective metrics of the GVA distribution. This indicates a more pronounced right-skewed shape with fatter tails. In other words, the distribution is more concentrated in the right tail which indicates positive effects on GFCF. These findings align with the core objective of the CF program, which is designed to stimulate investment and infrastructure development in less developed EU regions. Our results confirm that CF funds play a particularly strong role in boosting capital formation, with a frontloaded effect on investment. Additionally, this seems to affect the GVA over a longer horizon. Thus, these findings suggest that the increase in GVA occurs through the stimulation of investment.
Table 5: Summary statistics for the distribution of the regional effect on gross fixed capital formation
7 years 14 years Average
[…]
Skewness 1.16 0.15 0.51
Kurtosis 5.12 2.62 3.58
@@ 6.2.2 Geographical analysis of the heterogeneous level effects
This pattern reveals a new geographical division within Europe. In this case, Southern Europe, which consists of relatively wealthier regions, experiences weaker CF effects compared to Eastern Europe, where poorer regions benefit the most from such a policy. As such, the newest Member States that gained their EU status post-2000 are the ones that seem to reap the benefits of the Cohesion Fund assistance. A similar pattern though less pronounced is apparent when focusing on the effects on GFCF. The largest effects are again observed in Eastern Europe, while Southern Europe experiences milder impacts. It is interesting to note here that in the 14-year case, Spanish regions exhibit positive effects on investment.
Figure 6: Geographical Dispersion of GVA effects
Figure 7: Geographical Dispersion of GFCF effects
6.2.3 Heterogeneous impact of the Cohesion Fund on GVA growth
In this section we focus our attention on the regional growth effects of the Cohesion Fund expenditures. Specifically, Table 6 compares the observed GVA growth, ${g}^{y}_{t}$, from the year of initial treatment with its counterfactual GVA growth as estimated in Section 5.2.1. That is the growth that a region would have experienced had it not received the CF. The counterfactual GVA growth for the $t$-th year after treatment, $\widehat{g}^{y}_{t}$, is computed as:
$\widehat{g}^{y}_{t}=\left[\log(\widehat{\text{GVA}}_{t})-\log(\text{GVA}_{0})\right]\times 100$ (7)
where $\text{GVA}_{0}$ represents the GVA at the beginning of the programming period and $\widehat{\text{GVA}}_{t}$ is the estimated counterfactual GVA at the end of the $t$-th year after the treatment. [^13] Table C.1 in Appendix C repeats the same analysis, excluding outlier regions based on the pre-treatment RMSE To better assess the heterogeneous effects, we categorize the treated regions into five percentiles (column 1) based on their initial GVA per capita (column 2). The observed and counterfactual GVA growth rates, shown in columns four and five, respectively, are the average growth across regions of each percentile. We do this over two consecutive 7-year cycles, mimicking the actual 7-year budgeting programming periods of the EU.
Table 6: GVA per capita growth rates by years under treatment (group averages)
Percentile Initial GVA $g^{y}$, in % $\widehat{g}^{y}$, in % $g^{y}-\widehat{g}^{y}$
Effect After First 7-Year Treatment Period
0-10th 4137.21 32.7 11.2 21.5
10th-25th 5273.65 28.8 10.01 18.79
25th-50th 8416.783 24.9 12.55 12.35
50th-75th 11928.78 15.57 14.01 1.56
75th-90th 17153.59 11.58 15.67 -4.09
Effect After Second 7-Year Treatment Period
0-10th 4137.21 7.5 5.3 2.2
10th-25th 5273.65 9.4 11.3 -1.9
25th-50th 8416.783 8.09 5.57 2.52
50th-75th 11928.78 9.99 9.89 0.10
75th-90th 17153.59 11.89 13.56 -1.67
  • Notes: Initial GVA refers to the GVA per capita in the first year of the treatment. The last two columns depict the overall period growth rate for the cases in which the region is treated and the counterfactual, respectively.
Table 6 yields some interesting findings. The impact of EU’s Cohesion Fund is front-loaded, peaking within the first seven years of inclusion in the program. However, the effects exhibit significant heterogeneity, with a clear negative relationship between the initial GVA per capita and the estimated impact. That is, the regions that have benefited the most are those in the lower percentiles of the initial GVA distribution - specifically, those below the 50th percentile. Regions belonging in the 0-10th, 10th-25th, and 25th-50th percentiles experienced an overall GVA growth that was 21.5, 19 and 12 percentage points higher than their respective counterfactual scenario (i.e., in the absence of EU funding), respectively. In contrast, the benefits in terms of GVA growth diminish for the relatively wealthier regions, such as those in the 50th-75th percentile. For the highest-income regions, the effect turns to negative. This suggests that the Cohesion Fund may have had limited or even adverse effects in the wealthiest regions. When comparing across different funding cycles, we observe that the growth effects persist but diminish significantly over time. However, it is important to note that the second cycle, i.e., the 14 years under treatment, overlaps with the Global Financial Crisis, which may have affected the outcomes. Nonetheless, even during the Crisis period, it seems that the CF payments seem to have assisted a large number of recipient regions across the initial GVA distribution to attain an average GVA growth rate larger than the counterfactual case.
Table 7 is analogous to Table 6 but it computes overall investment growth. The counterfactual investment growth, $\widehat{g}_{t}^{I}$, is defined similarly to equation (7). For comparability with the above discussion, we categorize treated regions into the same five percentiles based on their initial GVA per capita. The results reinforce the discussion on the heterogeneous GVA effects discussed above. As with GVA growth, the impact of the EU’s Cohesion Fund on investment overall growth is frontloaded, peaking within the first seven years. The effect is also more pronounced in relatively poorer regions, with the higher impact estimated for the poorest 0-10th percentile, followed by the 10th-25th and 25th-50th percentiles. Compared to the effect on GVA growth, the effect on overall investment growth is significantly larger in magnitude. For instance, the poorest percentile experienced an investment growth rate that was 57.45% percentage points higher than in the counterfactual scenario (i.e., without EU funding). This impact is more than 2.5 larger than the estimated effect on GVA growth. This is intuitive as the CF is designed and targeted to stimulate investment and infrastructure in relatively poorest EU regions. Another key difference is that while the effects on GVA growth persist, the benefit to investment growth dissipate faster in the second-year cycle.
Table 7: Investment growth rates by years under treatment (group averages)
Percentile Initial GVA $\widehat{g}^{I}$, in % $g^{I}$, in % $\widehat{g}^{I}-g^{I}$
Effect After First 7-Year Treatment Period
0-10th 4137.21 66.67 9.22 57.45
10th-25th 5273.65 48.38 6.22 42.16
25th-50th 8416.783 25.99 7.74 18.25
50th-75th 11928.78 32.48 17.45 15.03
75th-90th 17153.59 32.87 21.47 11.40
Effect After Second 7-Year Treatment Period
0-10th 4137.21 -19.09 -11.14 -7.95
10th-25th 5273.65 -0.80 -5.26 4.46
25th-50th 8416.783 -7.47 -4.89 -2.58
50th-75th 11928.78 13.66 4.52 9.14
75th-90th 17153.59 25.71 12.68 13.03
  • Notes: Initial GVA refers to the GVA per capita in the first year of the treatment. The last two columns depict the overall period growth rate for the cases in which the region is treated and the counterfactual, respectively.
6.2.4 Analysis of the effects by country
Given that the Cohesion Fund is, by design, allocated to countries (rather than to regions, as with the rest of the EU’s Structural Funds) it is informative to examine the economic impact of the Cohesion Fund at a more aggregate level. In particular, given that we have estimated region- and time-specific treatment effects we can use equation (4) in order to obtain country-specific treatment effects (i.e. in this case the set $\mathcal{N}$ contains the NUTS2 regions of each of the eligible countries).
The following Tables 8 and  9 present the country-specific treatment effects for GVA per capita and investment, respectively, in the 7th and 14th year under treatment. We rank the countries in ascending order based on their initial GVA per capita, starting with the lowest initial GVA per capita, i.e., Bulgaria, and ending with the highest, i.e., Cyprus.
As expected, most of the results observed at the regional level are confirmed when aggregated at the country level. Specifically, the impact of the EU Cohesion Fund is frontloaded, with the largest effects estimated within the first seven-year cycle. Additionally, the magnitude of the impact is negatively correlated with a country’s initial GVA per capita, meaning poorer countries tend to experience stronger positive effects. These differences in GVA growth stem from the causal impact of regions within each country receiving Cohesion Fund support. The eight poorest countries in terms of initial GVA —Bulgaria, Romania, Lithuania, Hungary, Latvia, Poland, Slovakia, and Estonia—generally exhibit positive effects on both GVA and investment growth. As before, the effect on investment growth is more pronounced compared to the effect on GVA growth. Romania, Latvia, and Lithuania stand out with particularly high GVA growth rates over the first seven-year cycle, although some of these gains diminish over time. In contrast, Hungary and Slovakia display more moderate yet stable positive effects. The effect on investment growth is more volatile than GVA growth. While Bulgaria, Romania, Lithuania, and Latvia experience significant investment growth within the first seven years, their long-term effects tend to decline and/or even turning negative. Poland and Greece, in contrast, show more sustained investment effects over the two cycles. Cyprus experiences a dramatic investment decline from the first seven-year to the second seven-year cycle. Wealthier countries such as Spain and Portugal show weaker or even negative effects, suggesting diminishing returns to EU funding in economies with higher GVA per capita.
Table 8: Country-specific GVA growth effect, $\widehat{g}^{y}-g^{y}$
GVA pc Years under treatment
Country 7 14
Bulgaria (BG) 2991 14.1 3.36
Romania (RO) 4099 32.5 -2.17
Lithuania (LT) 5204 27.43 2.8
Hungary (HU) 5254 8.96 -4.29
Latvia (LV) 5300 30.93 2.52
Poland (PL) 5898 10.1 7.46
Slovakia (SK) 7352 22.44 7.81
Estonia (EE) 7775 16.42 -7.07
Croatia (HR) 9434 -6.25 3.24
Czech Republic (CZ) 9925 14.2 -0.61
Portugal (PT) 11954 -0.35 -5.47
Slovenia (SI) 12334 9.08 -2.92
Greece (EL) 12585 -1.26 8.14
Malta (MT) 13159 -4.64 9.47
Spain (ES) 15817 -4.94 -5.92
Cyprus (CY) 18192 0.12 -10.06
  • Notes: Cyprus, Estonia, Latvia, Malta constitute of only one region; as such, the average treatment effect is the same irrespective of the level of aggregation. $g^{y}$ denotes is the growth rate in the treated case and $\widehat{g}^{y}$ is the growth rate in the counterfactual scenario. The initial year of treatment differs across countries.
Table 9: Country-specific investment growth effect, $\widehat{g}^{I}-g^{I}$
GVA pc Years under treatment
Country 7 14
Bulgaria (BG) 2991 66.3 11.24
Romania (RO) 4099 66.98 -12.97
Lithuania (LT) 5204 61.63 -19.45
Hungary (HU) 5254 16.73 6.19
Latvia (LV) 5300 57.91 -23.62
Poland (PL) 5898 6.65 16.37
Slovakia (SK) 7352 9.78 4.02
Estonia (EE) 7775 65.93 -8.21
Croatia (HR) 9434 -0.06 2.11
Czech Republic (CZ) 9925 15.11 -0.51
Portugal (PT) 11954 17.08 -11.84
Slovenia (SI) 12334 15.92 -26.23
Greece (EL) 12585 19.26 12.62
Malta (MT) 13159 22.08 7.54
Spain (ES) 15817 3.63 15.65
Cyprus (CY) 18192 29.03 -49.71
  • Notes: See notes of Table 9.

### v2

@@ 6 Results
In this section we present the empirical estimates of the impact of Cohesion Fund payments on regional economic performance. In particular, we estimate region- and time-specific causal effects using the model specified in equation (3), including the employment-to-population (EMP) ratio, $X_{it}$, as a time-varying covariate. The EMP ratio is computed as the total number of employed individuals divided by the working-age population in each region at each time point. This variable is an important confounder to control for, as it is strongly correlated with GVA per capita and serves as a proxy for the strength of the regional labor market. Regions with higher EMP ratios typically exhibit stronger economic performance. Notably, treated regions in our sample tend to have significantly lower EMP ratios relative to the EU average, underscoring the need to adjust for this imbalance in the analysis. Following a similar argument, the employment-to-population ratio is included as a covariate in the estimation of the investment effects of the Cohesion Fund transfers. Specifically, as the employment-to-population ratio is known to exhibit a cyclical pattern and can serve as a proxy of the business cycle, its inclusion in the specification allows for a better identification of the variation in the outcome variable (in the case, the rate of investment). [^17] The estimation is conducted by using the gsynth library (Xu and Liu, 2021) in the statistical software R (R Core Team, 2021). Note that EMP is the only covariate available at the regional level for the whole timeframe available. This fact underscores the importance of accounting for unobserved councounders by use of interactive fixed effects, as in 3. Without allowing for these unobserved factors, the model at a regional level could be highly misspecified due to the lack of regional data.
The rest of the section is organized as follows: firstly, in section 6.1 we present the average treatment effect on the treated. Next, in section 6.2, we delve into the heterogeneous nature of the effects; specifically, section 6.2.1 examines the time-varying distribution of the effects, section 6.2.2 focuses on the spatial patterns that emerge and section 6.2.3 examines the relationship between the income level of the regions and the estimated effects. Section 6.3 examines the evolution of regional income heterogeneity over time by means of the Gini coefficient. Lastly, section 6.4 provides some graphical evidence on the non-linear nature of the relationship between funding intensity and its GVA effects.
@@ 6.1 Average Treatment Effect
We start by presenting the average treatment effect across all regions as estimated in section 5.2.1; specifically, see equation (4). Figures 4 and  4 present the average treatment effect on the y-axis against years relative to treatment for GVA per capita and Gross Fixed Capital Formation (GFCF), respectively. The vertical red line that crosses the x-axis in year 0 represents the initial year in which each region received CF payments. Note that this year might differ for each region as EU countries became eligible for the CF payments at different years and, as such, the regions in our sample received payments at different points in time. Thus, the treatment effect is shown by the black line for each year that follows the zero year. The grey shaded area represents the upper and lower bound of the $95\%$ confidence interval calculated using a parametric bootstrap technique as in Xu (2017) (see Appendix B for more details).
@@
Figures 4 and  4 show that on average CF payments have a persistent positive effect on both the GVA and GFCF of a region that receives them (treated) compared to the counterfactual scenario where these funds are not provided by the EU. [^18] Figure D.1 in Appendix D plots both the observed and the respective estimated counterfactual average GVA in logs as well as in growth rates. Their difference is the average treatment effect presented in Figure 4. Additionally, Figure D.2 depicts the same series for the growth rate of GVA.
The average treatment effects depicted on the y-axis of Figure 4 indicate that within the first seven years under treatment, Cohesion Fund transfers are associated with an increase of almost 17% in GVA per capita. By the end of three consecutive seven-years transfer cycles (i.e., after 21 years after inclusion in the CF program), GVA per capita is, on average, 30% higher than it would have been in the absence of CF (in log scale [^19] Suppose a hypothetical treated average EU region $i$ that received treatment in period 0. If $\widehat{ATT}_{t}=0.3$, then $\log\!\left(\frac{\text{GVA}_{it}}{\widehat{\text{GVA}}_{it}(0)}\right)=0.3$, so $\frac{\text{GVA}_{it}}{\widehat{\text{GVA}}_{it}(0)}=e^{0.3}\approx 1.35$, i.e. $\text{GVA}_{it}\approx 1.35\times\widehat{\text{GVA}}_{it}(0)$, where $\widehat{\text{GVA}}_{it}(0)$ is GVA in period $t$ under the counterfactual scenario that region $i$ had not received CF. As such, the majority of the income increase appears to materialize in the early years of the implementation of the policy programme, while further gains occur but are relatively smaller. This is an indication that the CF programme effects are persistent but quite frontloaded. In terms of the average impact on investment (GFCF), as depicted in Figure 4, we observe again a persistent positive impact that appears to be frontloaded; however, the average impact is both larger in magnitude and exhibits a more volatile path over time. Specifically, within the first seven years under treatment, the share of investment is almost 45% higher. Following this, the impact plateaus and exhibits a slight decline; nonetheless, by the end of three consecutive seven-year periods, the share of investment is almost 40% higher.
Using the estimates presented in Figure 4, we compute the net benefit for the average treated region due to its inclusion in the CF expressed in levels of GVA per capita. Specifically, GVA per capita increases by €815 more that it would under the counterfactual scenario of no Cohesion Fund support. [^20] Specifically, for the hypothetical EU27 treated region, GVA per capita rises by €2168 over the first seven years of treatment—from €9,064 in the first year to €11,232 in year seven. Under the counterfactual scenario, the corresponding increase is €1353 —from €8618 to €9971 in year seven. Similarly, on the 27th year under treatment the Cohesion Funds transfers lead to an increase in GVA per capita equal to 31.3%, while for the entire period the average increase is 28.5%. This implies that the hypothetical treated region experiences an increase in income that is €2961 larger than it would have been under the counterfactual scenario. Following a similar calculation, a hypothetical treated region experiences an increase in investment per capita that is €395 and €680 more than it would have been under the counterfactual scenario after 7 and 21 years of treatment, respectively. [^21] Figures C.2 and LABEL:fig:GFCF_MCNNvsTWFE in Appendix Comparison with TWFE show that the estimated average treatment effect of this section is quantitatively very close to the analogous estimate obtained from the traditional Two-Way Fixed Effects approach.
@@
Note: Average Treatment Effect on the Treated ($ATT$) over time (black solid line) and $95\%$ confidence intervals (gray shaded area) for the log of GVA per capita (left panel) and the log of GFCF/GVA (right panel).
@@ 6.2.1 Quantitative analysis of the heterogeneous effects
The previous section focused on the average treatment effect across all regions that have received CF payments. However, a key advantage of our empirical strategy is that it allows us to explore the heterogeneous effects of CF across the different regions in our sample.
Figure 5 illustrates the distribution of treatment effects on log real GVA per capita across treated regions at selected years relative to treatment. The kernel density estimates indicate that the distribution is centered near zero one year prior to treatment, as expected (see the bottom panel of Figure 5). Post-treatment, the distribution gradually shifts rightward, with increasing mass above zero at years 7, 14, and 21, suggesting that the policy intervention generated increasingly positive economic effects over time. The widening of the distribution at longer horizons also indicates growing heterogeneity in regional responses.
Figure 6 shows analogous results for log real gross fixed capital formation (GFCF). Again, the pre-treatment distribution is centered near zero, while post-treatment distributions, particularly at years 14 and 21, shift notably to the right and exhibit longer right tails. This pattern suggests that investment responses to the treatment are not only positive on average but also more skewed, indicating that some regions experienced particularly large in magnitude increases in investment.
The analysis in this section indicates that focusing solely on average treatment effects (as in section 6.1) obscures substantial heterogeneity in the effects of EU Cohesion Funds (CF). Recognizing this heterogeneity can provide valuable insights for EU policymakers when determining the optimal allocation of EU resources.  [^22] Indicatively, the observed heterogeneity could be related to the intensity of the treatment, i.e. the size of the Cohesion Fund expenditures relative to GVA. If a maximum desired intensity is identified, this provides flexibility to reallocate the resources and achieve better outcomes. This is a novelty of our approach, as most existing papers usually limit their analysis to the estimation of an average effect across regions, ignoring any potential heterogeneity. We touch upon this issue in more detail in section 6.4.
@@
Table 3 presents summary statistics for the time-varying distributions of the CF effects on regional gross value added (GVA) and gross fixed capital formation (GFCF), corresponding to the distributions shown in Figures 5 and  6, respectively. Starting with the effects on GVA, in the top panel of the Table, the results indicate that participation in the CF program tends to have a predominantly positive impact on GVA for the majority of treated regions (see first panel). Specifically, during the first seven-year cycle, $75.32\%$ of treated regions experience positive effects; this increases to $80.2\%$ after 14 years and marginally declines to $78.2\%$ after 21 years. On average, $80\%$ of the treated regions benefit from the CF in terms of GVA per capita. The skewness of the distribution (second row) shows a positively skewed effect in the first and second cycles, indicating that while most regions experience modest gains, a smaller number benefit disproportionately. However, skewness declines by the third cycle, suggesting that the most extreme positive effects become less frequent over time. Kurtosis follows a similar trajectory, although relatively moderate in earlier cycles, the average value rises to 1.99, implying a distribution with heavier tails—i.e., some regions experience particularly large or small effects relative to the mean.
The second panel reports the distributional characteristics of CF effects on GFCF (i.e., from Figure 6). Although most of the patterns observed for GVA also hold here, notable differences emerge. The impact on regional investment is significantly larger in magnitude and affects a greater proportion of treated regions. Nearly $89\%$ of regions experience a positive effect on investment in the first seven years—substantially higher than the corresponding figure for GVA. This share increases to almost $91\%$ after 14 years and then declines to $74\%$ after 21 years, with an overall average of $88.5\%$. The distribution of the investment effects also differs in shape. Skewness is higher in the early years and even turns slightly negative by the third cycle, while kurtosis consistently exceeds that of the GVA distribution. This suggests a more pronounced right-skewed shape early on, with fatter tails, meaning that although most regions benefit, some see particularly strong gains in investment. These findings align with the core objective of the CF program, which is to stimulate investment and infrastructure development in less developed EU regions. The frontloaded nature of these effects on GFCF implies that capital formation is an early outcome of CF participation. Moreover, this investment may serve as a transmission channel through which the observed longer-term gains in GVA are realized.
Table 3: Summary statistics for the distribution of the regional effects
@@
GVA per capita
$\%$ of regions with $\widehat{\tau}_{i}>0$ 75.32 80.2 78.2 80.2
Skewness 0.17 0.18 -0.16 0.04
Kurtosis 2.74 1.87 2.19 1.99
GFCF
$\%$ of regions with $\widehat{\tau}_{i}>0$ 88.5 90.6 73.9 88.5
Skewness 0.97 0.39 -0.04 0.84
Kurtosis 4.31 2.56 2.25 3.78
  • Notes: In the 21st year under treatment the sample consists of 92 regions, as the regions in Croatia are not treated.
  • $\dagger$ This column depicts the average based on the entire post-treatment period.
@@ 6.2.2 Geographical analysis of the heterogeneous level effects
This pattern reveals a new geographical division within Europe. In this case, Southern Europe, which consists of relatively wealthier regions, experiences weaker CF effects compared to Eastern Europe, where poorer regions benefit the most from such a policy. As such, the newest Member States that gained their EU status post-2000 are the ones that seem to reap the benefits of the Cohesion Fund assistance.
A similar pattern emerges when focusing on the effects on investment. As can be seen in Figure 8 the largest effects are again observed in Eastern Europe, while Southern Europe experiences milder impacts.
Figure 7: Spatial distribution of GVA per capita effects
Figure 8: Spatial distribution of Investement effects
6.2.3 Heterogeneous impact of the Cohesion Fund by income group
In this section we examine a different aspect of the observed heterogeneity in the estimated treatment effects. Particularly, we examine whether the effects depend on the relative position of a region across the income distribution, that is, whether poorer/richer regions experience different effects from the cohesion transfers. To do so, we categorize the treated regions into five quintiles based on their GVA per capita in the last year before the treatment (i.e., receiving CF payments for the first time). Figure 9 depicts $\widehat{ATT}_{t}$, separately for the five groups of regions. More precisely, the horizontal axis shows years relative to the start of the CF program (year 0), while the vertical axis measures the $\widehat{ATT}_{t}$. Each line represents a different quintile of regions ranging from the poorest $20\%(0\textendash 20\%)$ to the richest $20\%(80\textendash 100\%)$.
The graph shows substantial heterogeneity in the impact of the Cohesion Fund program across these quintiles. Regions in the lowest quintile (see red line) exhibit the largest $\widehat{ATT}_{t}$ in the years following treatment, while those in the highest quintile consistently display the smallest effects (see orange line). This pattern suggests that the program was significantly more beneficial for poorest regions, with treatment effects increasing over time and remaining persistently higher compared to regions in the highest quintiles. The widening gap in the estimated $\widehat{ATT}_{t}$ between the lowest and highest quintiles over the post-treatment period implies that the policy program contributed to the strengthening of the relative position of the poorest regions, thus ameliorating the cross-region income disparities.
Figure 9: $\widehat{ATT}$ by quintile based on pre-treatment GVA per capita distribution.
Figure 10 zooms in to examine the actual and counterfactual trajectories of log GVA per capita over time for each group of regions. The panels again group regions into quintiles based on their GVA per capita in the last pre-treatment year, from the poorest $(0\textendash 20\%)$ to the richest $(80\textendash 100\%)$. Within each group, the graph plots two series: the observed evolution of log real GVA per capita and its counterfactual (i.e., what GVA per capita would have been in the absence of the CF program). The difference between the observed (red solid) and counterfactual (light blue dashed) lines reflects the estimated treatment effect for each group. Consistent with the previous figure, the poorest regions (those in the $(0\textendash 20\%)$ and $(20\textendash 40\%)$ quintiles) display a clear and widening gap between observed and counterfactual GVA per capita in the years following treatment, indicating that they experienced substantial and sustained gains. In contrast, for richer regions (particularly those in the $(80\textendash 100\%)$ quintile), the two trajectories remain nearly parallel, suggesting minimal impact from the program. Interestingly, for the regions in the $(60\textendash 80\%)$ quintile the two trajectories coincide by the end of the period, indicating that the impact dissipates. This disaggregated analysis reinforces the earlier conclusion that the CF program had strong, persistent effects in the regions in the lower end of the income distribution. By enabling poorer regions to grow faster than their projected counterfactual paths, the program appears to have contributed significantly to reducing income disparities within the EU. Moreover, the quantitatively smaller effect estimated in richer regions suggests that the program may have limited capacity to generate impact the regions where the economic fundamentals appear relatively stronger. This could indicate a need for a more tailored design of the policy program, including a potential redistribution of funds from richer to poorer regions.
Figure 10: $\widehat{ATT}$ by quintile based on pre-treatment GVA per capita distribution.
A similar picture emerges when we shift our focus to the evolution of investment following the disbursement of the Cohesion Fund payments. As above, Figure 11 depicts the observed and the counterfactual time-path of the investment to GVA ratio in the log-scale by income quintile. As can be gleaned from the Figure, the regions belonging to the three lowest quintiles experience the largest benefits in terms of investment with the gap between observed and counterfactual widening over time (most notably in the ‘poorest of the poor’ regions of the lowest quintile). On the contrary, in the two highest quintiles composed of the richer regions one observes that the difference between the two trajectories is rather flat, with the gap disappearing by the end of the period under consideration in $60-80\%$ percentile. This explains the negligible effect of the CF transfers in terms of GVA per capita for this group that was analyzed above.
Figure 11: $\widehat{ATT}$ by quintile based on the distribution of the pre-treatment investment to GVA ratio.
6.3 Regional heterogeneity over time
As shown in the previous sections, the Cohesion Fund exerts a persistent positive impact on the output of the ‘poorest of the poor’ regions, that is, regions belonging to the bottom quintiles of the income distribution. Given these persistent effects, it is of particular interest to examine how income disparities across the EU regions have evolved over time by comparing the observed outcomes with the counterfactual scenario in which no region receives the relevant transfers. To this end, we employ one of the most commonly used statistical measures for income inequality, namely, the Gini coefficient. Specifically, we compute the time path of the Gini coefficient for both the observed (blue-dotted line) and counterfactual (red dashed-dotted line) levels of GVA per capita across all EU regions from 1995 to 2022. For the non-treated regions, the observed time series are used in both cases. We refer to this measure as the regional Gini coefficient, as it captures the cross-regional dispersion of GVA per capita over the 1995-2022 period. From a visual inspection of Figure 12 two interesting observations emerge. First, the observed Gini coefficient declines steadily from around 0.31 to below 0.28 by the end of sample period. In sharp contrast, under the counterfactual distribution, i.e., assuming the absence of the CF program, the dispersion of GVA across regions fluctuates around 0.31 and even slightly increases over time.
The decline begins in the early 2000s and coincides with the enlargement of the EU following the accession of the Central and Eastern European members which, as we discussed earlier, were the main recipients of the CF. The divergence between the counterfactual and actual regional Gini underlines the important role of the Cohesion Fund transfers in mitigating income disparities across the EU’s regions. The strong positive effects on the output of the new Member States contributed to reducing regional divergence relative to the counterfactual no-CF scenario.
Figure 12: Observed (black solid line) and counterfactual (blue dotted line) Gini index based on GVA per capita